At a certain speed, the root-mean-square-speed of the molecules of hydrogen in a sample 

of gas is 1055 ms-1

. Compute the root-mean square speed of molecules of oxygen at the 

same temperature.


1
Expert's answer
2022-04-11T17:04:10-0400

u(H2) = 1055 ms^-1;

M(H2) = 2 g/mol;

M(O2) = 32 g/mol;

T = u(H2)^2 * M(H2)/(3 * R),

u(O2) = (3 * R * T/M(O2))^1/2 = (3 * R * u(H2)^2 * M(H2)/(3 * R *M(O2)))^1/2 = (u(H2)^2 * M(H2)/M(O2))^1/2 = (1055^2 * 2/32)^1/2 = 69564^1/2 = 263.7 ms^_1.

Answer: 263.7 ms^-1.